# Human vs. Agent in Task-Oriented Conversations

### v1

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region; December 7–10, 2025; Xi’an, China.booktitle: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (SIGIR-AP ’25), December 7–10, 2025, Xi’an, Chinaisbn: 979-8-4007-2218-9/2025/12.doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Information retrieval
@@ 1. Introduction
However, the effectiveness of simulation-driven approaches critically depends on the fidelity and reliability of the underlying LLM-simulated users. Researchers are increasingly employing LLM-based user simulations, yet the fundamental question of how closely these simulations replicate the behavioral characteristics of real human users remains inadequately addressed. While prior research (Zhu et al., 2024, 2025) has identified various limitations in user simulations—such as unnatural conversational patterns, limited contextual depth, or inconsistent personal adherence, particularly noted in conversational recommender systems (CRS)—a significant gap remains. Systematic, direct comparisons of LLM-simulated user behaviors against actual human user behaviors under rigorously controlled, near-identical conditions are notably scarce. Understanding the precise nature and magnitude of behavioral discrepancies is crucial for enhancing the realism and utility of simulations.
@@ 2.1. User Simulation
User simulation has emerged as a valuable technique across various research domains, enabling cost-effective testing and evaluation without requiring extensive data from real users. Before LLMs, user simulators were widely employed for evaluating recommender systems (Afzali et al., 2023; Zhang and Balog, 2020) and task-oriented conversational systems (Sun et al., 2021; Sekulić et al., 2022).
The advent of LLMs has significantly expanded simulation capabilities, revolutionizing multiple computational domains. These models excel in data augmentation by generating high-quality synthetic data (Askari et al., 2023; Huang et al., 2023; Whitehouse et al., 2023) and enable sophisticated social behavior modeling through multi-agent simulations (Park et al., 2023). Additionally, LLMs have proven valuable for automating conversational system evaluation (Chan et al., 2023; Hu et al., 2023) and enhancing recommender systems through agent-based approaches (Wang et al., 2023b; Zhang et al., 2024; Wang et al., 2023a; Zhu et al., 2024). These diverse applications demonstrate the versatility of LLM agents in addressing complex simulation challenges across various fields. Following these studies, we adopt the agent-based user simulation for our analysis.
@@ 2.2. Task-Oriented Conversational System
Task-oriented conversational systems aim to assist users in accomplishing specific objectives through natural dialogue, addressing real-world needs such as booking services, finding information, or completing transactions (Zhang et al., 2019; Peng et al., 2021). These systems have become increasingly prevalent in practical applications, including customer service, travel planning, restaurant reservations, and technical support, reflecting the diverse task-completion scenarios users encounter in daily life.
The emergence of LLMs has transformed approaches to building task-oriented conversational systems. While early work explored zero-shot applications of LLMs to these tasks (Hudeček and Dušek, 2023), subsequent research demonstrated that fine-tuning significantly improves performance (Hosseini-Asl et al., 2020; Gupta et al., 2022). More recent advances include instruction-tuning LLMs for specific subtasks, such as API argument filling (Mok et al., 2024), leveraging description-based approaches (Zhao et al., 2022), and developing sophisticated simulation frameworks (Luo et al., 2024; Sekulić et al., 2024). The automation of task-oriented chatbot development has also gained attention, with LLMs enabling more efficient creation of conversational flows (Sánchez Cuadrado et al., 2024).
@@
In game-theoretic settings, research (Xie et al., 2024) showed that LLM agents systematically differ from humans in trust-based interactions, typically exhibiting higher trust levels, while Fontana et al. (2025) discovered that LLMs behave more cooperatively than humans in the Prisoner’s Dilemma. Alsagheer et al. (2024) found that while LLMs show reasoning errors similar to humans, these similarities are superficial, and both groups respond differently to the same prompting strategies. Most recently, researchers have revealed that although LLMs display similar aggregate-level biases to humans (Horowitz and Plonsky, 2025), these patterns emerge from fundamentally different processes. LLMs exhibit strong recency biases and lack sophisticated phenomena, such as ”surprise triggers change”, which is observed in human behavior.
Under conversational settings, Huijzer and Hill (2023) demonstrated that LLMs exhibit behaviors remarkably similar to humans in controlled experiments. Jiaqi et al. (2025) highlighted that current human-LLM comparisons often rely on surface-level language similarities rather than deeper behavioral understanding.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region; December 7–10, 2025; Xi’an, Chinabooktitle: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (SIGIR-AP 2025), December 7–10, 2025, Xi’an, Chinadoi: 10.1145/3767695.3769490isbn: 979-8-4007-2218-9/2025/12ccs: Information systems Information retrieval
@@ 1. Introduction
However, the effectiveness of simulation-driven approaches critically depends on the fidelity and reliability of the underlying LLM-simulated users. Researchers are increasingly employing LLM-based user simulations, yet the fundamental question of how closely these simulations replicate the behavioral characteristics of real human users remains inadequately addressed. While prior research (zhu2024reliable; zhu2025llm) has identified various limitations in user simulations—such as unnatural conversational patterns, limited contextual depth, or inconsistent personal adherence, particularly noted in conversational recommender systems (CRS)—a significant gap remains. Systematic, direct comparisons of LLM-simulated user behaviors against actual human user behaviors under rigorously controlled, near-identical conditions are notably scarce. Understanding the precise nature and magnitude of behavioral discrepancies is crucial for enhancing the realism and utility of simulations.
@@ 2.1. User Simulation
User simulation has emerged as a valuable technique across various research domains, enabling cost-effective testing and evaluation without requiring extensive data from real users. Before LLMs, user simulators were widely employed for evaluating recommender systems (afzali2023usersimcrs; zhang2020evaluating) and task-oriented conversational systems (sun2021simulating; sekulic2022evaluating).
The advent of LLMs has significantly expanded simulation capabilities, revolutionizing multiple computational domains. These models excel in data augmentation by generating high-quality synthetic data (askari2023generating; huang2023chatgpt; whitehouse2023llm) and enable sophisticated social behavior modeling through multi-agent simulations (park2023generative). Additionally, LLMs have proven valuable for automating conversational system evaluation (chan2023chateval; hu2023unlocking) and enhancing recommender systems through agent-based approaches (wang2023recagent; zhang2024agentcf; wang2023rethinking; zhu2024reliable). These diverse applications demonstrate the versatility of LLM agents in addressing complex simulation challenges across various fields. Following these studies, we adopt the agent-based user simulation for our analysis.
@@ 2.2. Task-Oriented Conversational System
Task-oriented conversational systems aim to assist users in accomplishing specific objectives through natural dialogue, addressing real-world needs such as booking services, finding information, or completing transactions (zhang2019dialogpt; peng2021soloist). These systems have become increasingly prevalent in practical applications, including customer service, travel planning, restaurant reservations, and technical support, reflecting the diverse task-completion scenarios users encounter in daily life.
The emergence of LLMs has transformed approaches to building task-oriented conversational systems. While early work explored zero-shot applications of LLMs to these tasks (hudevcek2023llms), subsequent research demonstrated that fine-tuning significantly improves performance (hosseini2020simple; gupta2022instructdial). More recent advances include instruction-tuning LLMs for specific subtasks, such as API argument filling (mok2024llm), leveraging description-based approaches (zhao2022description), and developing sophisticated simulation frameworks (luo2024duetsim; sekulic2024reliable). The automation of task-oriented chatbot development has also gained attention, with LLMs enabling more efficient creation of conversational flows (sanchez2024automating).
@@
In game-theoretic settings, research (xie2024can) showed that LLM agents systematically differ from humans in trust-based interactions, typically exhibiting higher trust levels, while fontana2025nicer discovered that LLMs behave more cooperatively than humans in the Prisoner’s Dilemma. alsagheer2024comparing found that while LLMs show reasoning errors similar to humans, these similarities are superficial, and both groups respond differently to the same prompting strategies. Most recently, researchers have revealed that although LLMs display similar aggregate-level biases to humans (horowitz2025llm), these patterns emerge from fundamentally different processes. LLMs exhibit strong recency biases and lack sophisticated phenomena, such as ”surprise triggers change”, which is observed in human behavior.
Under conversational settings, huijzer2023large demonstrated that LLMs exhibit behaviors remarkably similar to humans in controlled experiments. jiaqi2025comparative highlighted that current human-LLM comparisons often rely on surface-level language similarities rather than deeper behavioral understanding.